OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (anonymous):

y+2=1(x-6) y=x-8